# Generative Diffusion Models for Resource Allocation in Wireless Networks

### v1

@@ IV GNN-Parametrizations for GDM Policies
GNNs process graph data through a cascade of $L$ graph convolutional network (GCN) layers. Inputs are node signals (features) and graph shift operators (GSO) while outputs are node embeddings. Each GCN layer $\boldsymbol{\Psi}^{(\ell)}$ is a nonlinear aggregation function obtained by the composition of a graph convolutional filter and a pointwise nonlinearity $\varphi$ (e.g., relu),
${\mathbf{Z}}^{(\ell)}\!=\!\boldsymbol{\Psi}^{(\ell)}\left({\mathbf{Z}}^{(\ell-1)};{\mathbf{H}},\Theta^{(\ell)}\right)\!=\!\varphi\left[\,\sum_{k=0}^{K}\;{\mathbf{H}}^{k}{\mathbf{Z}}^{(\ell-1)}\boldsymbol{\Theta}_{k}^{(\ell)}\right]\!.$ (9)
In (9), $\Theta^{(\ell)}=\{\boldsymbol{\Theta}_{k}^{(\ell)}\in{\mathbb{R}}^{F_{\ell-1}\times F_{\ell}}\}_{k=0}^{K}$ is a set of learnable weights, and ${\mathbf{Z}}^{(\ell-1)}\in{\mathbb{R}}^{N\times F_{\ell-1}}$ is the input node signal to layer $\ell$. The GSO, ${\mathbf{H}}$, encodes the underlying connectivity of the network, which is the network state in our case.
For improved training, we take advantage of normalization layers and residual connections. To this end, we redefine $\varphi$ in (9) as the composition of a normalization layer followed by a pointwise nonlinearity while the first term in the sum, ${\mathbf{Z}}\boldsymbol{\Theta}_{0}$, inherently represents a learnable residual connection.
We view ${\mathbf{x}}_{t}$ and ${\mathbf{t}}=t\mathbf{1}_{N}$ as node signals and introduce a read-in layer $\boldsymbol{\Phi}^{(0)}=(\boldsymbol{\Phi}_{{\mathbf{x}}},\boldsymbol{\Phi}_{{\mathbf{t}}})$ that adds sinusoidal-time embeddings to the input node features. That is, we have
${\mathbf{Z}}^{(0)}=\boldsymbol{\Phi}^{(0)}({\mathbf{x}}_{t},{\mathbf{t}})=\boldsymbol{\Phi}_{{\mathbf{x}}}({\mathbf{x}}_{t})+\boldsymbol{\Phi}_{{\mathbf{t}}}({\mathbf{t}}),$ (10)
where $\boldsymbol{\Phi}_{{\mathbf{x}}}:{\mathbb{R}}^{N}\mapsto{\mathbb{R}}^{N\times F_{0}}$ is a multi-layer-perceptron (MLP) layer, and $\boldsymbol{\Phi}_{{\mathbf{t}}}:{\mathbb{R}}^{N}\mapsto{\mathbb{R}}^{N\times F_{0}}$ is a cascade of a sinusoidal time embedding and MLP layers. Finally, we add a readout MLP layer $\boldsymbol{\Phi}^{(L)}:{\mathbb{R}}^{N\times F_{L}}\mapsto{\mathbb{R}}^{N}$ that learns to predict the noise $\boldsymbol{\epsilon}$ from the output node embeddings.
@@ V Case Study: Power Control in Multi-User Interference Networks
Figure 1: Two dimensional slices of expert policy (blue) and GDM policy (orange) samples (normalized transmit power allocations) are shown for four transmitter-receiver pairs (nodes) from two example test networks. E.g., normalized transmit powers ${\mathbf{x}}_{1}/P_{\max}$ and ${\mathbf{x}}_{2}/P_{\max}$ for the first two nodes sampled from graphs ${\mathbf{H}}_{1}$ and ${\mathbf{H}}_{2}$ are plotted in the first column.
Figure 2: Comparison of the expert policy, GDM policy and baselines in terms of the ergodic mean-rate objective $f_{0}$ averaged across all networks and several percentiles of constraints ${\mathbf{f}}$ given by the ergodic rates of all receivers. A minimum rate requirement $f_{\min}$ is shown with a dashed, red line.
We consider the problem of power control in $N$-user interference channels. A similar setup has been investigated in [16] and should be referred to for more details.
@@ V-A Wireless Network & Power Control Setup
To summarize the setup briefly, each network is comprised of transmitter-receiver pairs where each transmitter communicates only with its corresponding receiver while causing interference to other neighboring receivers. All network realizations are sampled from a given network configuration with $N=200$ transmitters-receiver pairs (nodes in our graphs) and a density of 8 pairs/km2. Following [16], we first drop the transmitters randomly in a square grid-world, and then drop the receivers in a certain neighborhood of their associated transmitters.
We optimize the transmit power levels ${\mathbf{x}}\in[0,P_{\max}]^{N}$ where $P_{\max}=10$ mW is the maximum transmit power budget of each transmitter. The channel bandwidth and noise power spectral density (PSD) are set to $W=20$ MHz and $N_{0}=-174$ dBm/Hz, respectively.
The network state ${\mathbf{H}}$ is the matrix of long-term channel gains which follow a log-normal shadowing with standard deviation of $7$ plus the standard dual-slope path-loss model. For a given ${\mathbf{H}}$, the short-term (instantaneous) channel gains $\widetilde{{\mathbf{H}}}$ vary following Rayleigh fading. To evaluate the performance of a policy ${\mathbf{x}}$, we define the instantaneous rate of receiver $i$ as
@@
where $x_{i}$ is the $i$th element in ${\mathbf{x}}$ and $\widetilde{h}_{ji}$ is the $(j,i)$th element in matrix $\widetilde{\mathbf{H}}$. Observe that a policy $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}})$ is determined only by the long-term gains, whereas the instantaneous rate depends on the short-term channel gains. Ergodic rates are defined as
${\mathbf{r}}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}}):=\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),\;\widetilde{{\mathbf{H}}}|{\mathbf{H}}}\left[\widetilde{{\mathbf{r}}}\big{(}{\mathbf{x}},\widetilde{{\mathbf{H}}}\big{)}\right].$ (12)
In (12), we take an expectation over the policy and the fading jointly. In our experiments, we evaluate the joint expectation over 200 timesteps, with each timestep spanning 10 ms. The utility constraints ${\mathbf{f}}$ impose a minimum ergodic rate requirement of $f_{\min}=0.5$ bps/Hz per receiver. The utility objective $f_{0}$ is the ergodic mean-rate given by $\mathbf{1}^{\top}{\mathbf{r}}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}})/N$.
V-B Implementation & Training Details for GDM Policy
We train the GNN-parametrized GDM policy over graphs whose edge weights are set to the normalized log-transformed long-term channel gains, i.e., $\log(1+h_{ij})$, where $h_{ij}$ denotes the co-channel and interference channel gains of $j$th receiver for $i=j$ and $i\neq j$, respectively. Our training dataset consists of $|{\mathcal{D}}_{{\mathbf{H}}}|=32$ random network state realizations of a given network configuration. The physical locations of the transmitter-receiver pairs, and hence the path-loss gains, remain fixed, whereas the shadowing loss varies across network realizations.
The GNN has $L=6$ layers, each with $F_{\ell}=1024$ hidden features and $K_{\ell}=2$ filter taps. We set the number of diffusion timesteps to $T=500$, use a cosine noise schedule $\beta_{t}$ and train the GDM policy to minimize the DDPM objective in (8) with a log-SNR weighting function $\omega(t)$. The training is executed for 1000 epochs with an ADAM optimizer, mini-batches of $1$ graph and $200$ graph signals, and a learning rate of $10^{-2}$.
@@ V-B Implementation & Training Details for GDM Policy
We apply an affine transform $[0,P_{\max}]\mapsto[-1/2,1/2]$ to map the policy space to a centered diffusion space. To sample from the diffusion model, we run (7) with standardized variables, invert the affine transform and constrain the generated policy samples to the support $[0,P_{\max}]^{N}$.
V-C (State-Augmented) Primal-Dual Expert Policy & Baselines
To generate samples from an expert policy, we first train a GNN-parametrized model via a state-augmented primal-dual (SA-PD) learning algorithm [16] and then sample trajectories from the dual dynamics of the SA-PD algorithm. We compare the expert and GDM policies with a full-power transmission baseline (FP) and a weighted minimum mean-square error (WMMSE) approach that maximizes the ergodic mean-rate.
@@ V-D Performance of GNN-Parametrized GDM Policy
Figure 3: Comparison of ergodic rates attained for example nodes sampled uniformly from a test network. Nodes are arranged in a decreasing order of ergodic rates under expert policy and red dashed line indicates the minimum rate requirement $f_{\min}$.
We evaluate the GDM policy on a test dataset of $|{\mathcal{T}}_{{\mathbf{H}}}|=8$ network states, sampled from the same family of networks as the training dataset $\mathcal{D}_{{\mathbf{H}}}$. Fig. 1 showcases several two-dimensional slices of learned $200$-dimensional solutions. We plotted the same node pairs across two different graphs in each column to highlight the variability of expert policy across different network states. We observe that expert policies are stochastic and usually multi-modal (e.g., top left plot), and the GDM policy matches the conditional expert distributions closely in general.
Various percentiles of ergodic rates and average ergodic rates are reported in Fig. 2. The expert policy solves the constrained power allocation problem with almost all constraints satisfied, while neither of the baselines does. The GDM policy performance mirrors that of the expert policy for average and percentile rates except for the lowest percentiles. The feasibility gap can be attributed to the supervised training loss not accounting directly for the sensitivity of the constraints.
In Fig. 3, we present the ergodic rates attained by a subset of transmitter-receiver pairs (nodes) of an example test network under the expert policy and the GDM policy. We note that the two policies agree for the majority of nodes and are strictly feasible. However, large deviations are not rare (e.g., nodes 5 and 11) and the GDM policy is not exempt from hallucinating spurious samples or limited coverage of the expert conditional distributions (e.g., bottom plots in Fig. 1). These are challenges conditional diffusion models are known to face.
@@ VI Conclusion
This work demonstrated that generative diffusion processes can imitate expert policies that sample from optimal solution distributions of stochastic network optimization problems. We employed a GNN to condition the generative process on a family of wireless network graphs. More broadly, we anticipate that our attempt at generative diffusion-based sampling of random graph signals will be of interest beyond resource optimization in wireless networks.
The feasibility gap and overall performance of the GDM policies can be improved by a training loss and generative process guided by the constraints. We leave constraint-aware, unsupervised training of GDM policies across a wider range of network topologies as future research directions.

### v3

@@ IV GNN-Parametrizations for GDM Policies
GNNs process graph data through a cascade of $L$ graph convolutional network (GCN) layers [32]. Inputs are node signals (features) and graph shift operators (GSO) while outputs are node embeddings. Each GCN layer $\boldsymbol{\Psi}^{(\ell)}$ is a nonlinear aggregation function obtained by the composition of a graph convolutional filter and a pointwise nonlinearity $\varphi$ (e.g., relu),
$\hskip-6.47491pt{\mathbf{Z}}^{(\ell)}\!=\!\boldsymbol{\Psi}^{(\ell)}\left({\mathbf{Z}}^{(\ell-1)};{\mathbf{H}},\Theta^{(\ell)}\right)\!=\!\varphi\left[\,\sum_{m=0}^{M_{\ell}}\;{\mathbf{H}}^{m}{\mathbf{Z}}^{(\ell-1)}\boldsymbol{\Theta}_{m}^{(\ell)}\right]\!.$ (9)
In (9), $\Theta^{(\ell)}=\{\boldsymbol{\Theta}_{k}^{(\ell)}\in{\mathbb{R}}^{F_{\ell-1}\times F_{\ell}}\}_{k=0}^{K}$ is a set of learnable weights, $M_{\ell}$ denotes the number of hops, and ${\mathbf{Z}}^{(\ell-1)}\in{\mathbb{R}}^{N\times F_{\ell-1}}$ is the input node signal to layer $\ell$. The GSO, $\mathbf{H}$}, encodes the underlying connectivity of the network, which is the network state in our case.
For improved and more stable training, we take advantage of normalization layers and residual connections. To this end, we redefine $\varphi$ in (9) as the composition of a normalization layer followed by a pointwise nonlinearity, while the first term in the sum, ${\mathbf{Z}}\boldsymbol{\Theta}_{0}$, inherently represents a learnable residual connection.
We view ${\mathbf{x}}_{k}$ and ${\mathbf{k}}=k\mathbf{1}_{N}$ as node signals and introduce a read-in layer $\boldsymbol{\Phi}^{(0)}=(\boldsymbol{\Phi}_{{\mathbf{x}}},\boldsymbol{\Phi}_{{\mathbf{t}}})$ that adds sinusoidal-time embeddings to the input node features. That is, we have
${\mathbf{Z}}^{(0)}=\boldsymbol{\Phi}^{(0)}({\mathbf{x}}_{k},{\mathbf{k}})=\boldsymbol{\Phi}_{{\mathbf{x}}}({\mathbf{x}}_{k})+\boldsymbol{\Phi}_{{\mathbf{k}}}({\mathbf{k}}),$ (10)
where $\boldsymbol{\Phi}_{{\mathbf{x}}}:{\mathbb{R}}^{N}\mapsto{\mathbb{R}}^{N\times F_{0}}$ is a multilayer perceptron (MLP) layer, and $\boldsymbol{\Phi}_{{\mathbf{t}}}:{\mathbb{R}}^{N}\mapsto{\mathbb{R}}^{N\times F_{0}}$ is a cascade of a sinusoidal time embedding and MLP layers. Finally, we add a readout MLP layer $\boldsymbol{\Phi}^{(L)}:{\mathbb{R}}^{N\times F_{L}}\mapsto{\mathbb{R}}^{N}$ that learns to predict the noise $\boldsymbol{\epsilon}$ from the output node embeddings.
@@ V Case Study: Power Control in Multi-User Interference Networks
Figure 1: Comparison of the test performance of GDM policy with the expert policy (SA) and other baselines. Leftmost and middle plots show the time evolution of the average and 5th percentile of ergodic rates, respectively. The minimum rate requirement $f_{\min}$ is shown with a dashed, red line. Rightmost plot shows the distribution of ergodic rates and constraint satisfaction percentages, evaluated up to $\tau=20$ and $\tau=200$ time steps. Solid black lines are drawn at the median values while the dashed black lines are for the 5th and 95th percentile values.
We consider the problem of power control in $N$-user interference channels. Similar setups have been investigated in [22, 23] and should be referred to for more details.
@@ V-A Wireless Network & Power Control Setup
To summarize the setup briefly, all network realizations are sampled from a family of network configurations with $N=100$ transmitters-receiver (tx-rx) pairs (also nodes in our graphs) and an average density of $12$ tx-rx pairs/ $\text{\,}{\mathrm{km}}^{2}$. For each network, we first drop the transmitters randomly in a square grid world, and each transmitter is paired with a neighboring receiver. Signals coming from all but their respective transmitters are treated as interference by the receivers.
We optimize the transmit power levels ${\mathbf{x}}\in[0,P_{\max}]^{N}$ where $P_{\max}=10$ mW is the maximum transmit power budget. The channel bandwidth and noise power spectral density (PSD) are set to $W=20$ MHz and $N_{0}=-174$ dBm/Hz, respectively. The network state $\mathbf{H}$} is the matrix of long-term channel gains which follow a log-normal shadowing with a standard deviation of $7$ plus the standard dual-slope path-loss model. For a given $\mathbf{H}$}, the short-term (instantaneous) channel gains $\widetilde{{\mathbf{H}}}$ vary following Rayleigh fading. To evaluate the performance of a policy $\mathbf{x}$}, we define the instantaneous rate of receiver $i$ as
@@
where $x_{i}$ is the $i$th component of $\mathbf{x}$} and $\widetilde{h}_{ji}$ is the $(j,i)$th element in matrix $\widetilde{\mathbf{H}}$. Observe that a policy $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}})$ is determined only by the long-term gains, whereas the instantaneous rate depends on the short-term channel gains. Ergodic rates are defined as
[…]
In (12), we take an expectation over the policy and the fading jointly. In our experiments, we draw $200$ samples from the trained GDM policy for each network and evaluate the joint expectation over 200 time steps, with each time step spanning 10 ms.
A minimum ergodic rate requirement of $f_{\min}=0.6$ bps/Hz is imposed for all receivers by setting the utility constraints ${\mathbf{f}}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}}):={\mathbf{r}}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}})-\mathbf{1}_{N}f_{\min}$ whereas the utility objective is the network-wide average of the ergodic rates given by $f_{0}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}}):=\mathbf{1}^{\top}_{N}{\mathbf{r}}(\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}}),{\mathbf{H}})/N$.
V-B State-Augmented Primal-Dual Expert Policy & Baselines
To generate samples from an optimal solution distribution, we first train a GNN-parametrized model via a state-augmented primal-dual (SA) learning algorithm as in [23]. The trained model is executed online for each given network $\mathbf{H}$} over a sufficiently long time window to generate trajectories of primal and dual iterates with near-optimality and feasibility guarantees. We collect the resulting primal iterates $\{{\mathbf{x}}^{\dagger}_{b}({\mathbf{H}})\}_{b=0}^{B-1}$, i.e., resource allocation vectors, in a buffer with capacity $B=500$. During GDM policy training, expert policy samples are uniformly drawn from the buffer, i.e., we have $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})=\mathrm{Uniform}\big[\{{\mathbf{x}}^{\dagger}_{0}({\mathbf{H}}),\ldots,{\mathbf{x}}^{\dagger}_{B-1}({\mathbf{H}})\}\big]$ for all ${\mathbf{H}}\sim\mathcal{D}_{{\mathbf{H}}}$. Note that while we parametrize the expert policy and run a state-augmented training algorithm over a training dataset of networks $\mathcal{D}_{{\mathbf{H}}}$, one can alternatively run the dual descent without the parametrization and store the resource allocation iterates for each instance ${\mathbf{H}}\sim\mathcal{D}_{{\mathbf{H}}}$.
We compare the GDM and expert policies with two baseline deterministic policies:
  1. 1.
    Average-power transmission policy: For each given $\mathbf{H}$}, we compute the time-average of the expert policy samples–similar to primal averaging in the optimization literature–and fix it, i.e., we set ${\mathbf{x}}({\mathbf{H}})\approx\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})}[{\mathbf{x}}^{\dagger}({\mathbf{H}})]$ at all time steps.
  2. 2.
    Full-power transmission policy (FP): All transmitters use all the transmission power available at all time steps, i.e., ${\mathbf{x}}({\mathbf{H}})=P_{\max}\mathbf{1}_{N}$.
V-C Implementation & Training Details for GDM Policy
For the GNN-parametrizations, the GSO representation of the network state $\mathbf{H}$} is a fully connected graph where nodes correspond to transmitter-receiver pairs, and the edge weights between nodes $i$ and $j$, $e_{ij}$, are set to log-normalized long-term channel gains given by $e_{ij}\propto\log_{2}\left(1+\frac{P_{\max}|h_{ij}|^{2}}{WN_{0}}\right)$. The GNN has $L=6$ layers, each with $F_{\ell}=128$ hidden features and $M_{\ell}=2$ hops (filter taps). We set the number of diffusion time steps to $K=500$, use a cosine noise schedule $\beta_{k}$ [33] and train the GDM policy to minimize the DDPM objective in (8) with a log-SNR weighting function $\omega(k)=\log\big(\mathrm{SNR}(k)\big)$, where $\mathrm{SNR}(k):=\alpha^{2}_{k}/\sigma^{2}_{k}$.
To evaluate our method, we draw a total of $128$ network realizations. Following a $5:1:2$ split, we obtain training, validation, and test datasets of size $|\mathcal{D}_{{\mathbf{H}}}|=80$, $|{\mathcal{V}}_{{\mathbf{H}}}|=16$, and $|{\mathcal{T}}_{{\mathbf{H}}}|=32$ networks, respectively. The expert policy solutions are obtained across all three datasets for evaluation and benchmarking purposes. We train the GDM policy over the training dataset $\mathcal{D}_{{\mathbf{H}}}$ for a maximum of $10^{4}$ epochs with an ADAM optimizer [34], an initial learning rate of $10^{-2}$, and a learning rate schedule that follows a cosine decay with warm restarts. In each epoch, we iterate over the whole dataset with mini-batches of $16$ graphs and sample $250$ graph signals, i.e., expert policy data ${\mathbf{x}}({\mathbf{H}})\sim\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, for each graph in the batch. Every $200$ epochs, we evaluate the checkpointed GDM model on the validation dataset ${\mathcal{V}}_{{\mathbf{H}}}$ and save the best model in terms of $5$th percentile ergodic rates. We test the saved model on ${\mathcal{T}}_{{\mathbf{H}}}$.
@@ V-C Implementation & Training Details for GDM Policy
We apply an affine transform $0,P_{\max}]\mapsto[-1/2,1/2]$ to map the policy space to a centered diffusion space. To sample from the diffusion model, we run the DDPM sampling equation (7) with standardized variables, invert the affine transform, and project the generated policy samples to the support $0,P_{\max}]^{N}$. [^1] During inference, we observed negligible difference in the quality of generations when we swapped the DDPM sampler with other samplers, e.g., a DDIM sampler or its accelerated counterpart with fewer denoising time steps.
Figure 2: Example two dimensional slices of expert policy (blue) and GDM policy (orange) samples are shown for two pairs of neighboring nodes from the test dataset. Transmit powers are normalized by $P_{\max}$.
@@ V-D Performance of GNN-Parametrized GDM Policy
[…]
In Fig. 2, example two-dimensional (2D) slices of $100$-dimensional learned GDM policies are overlaid with expert policy samples. The GDM policy generalizes to unseen test networks drawn from ${\mathcal{T}}_{{\mathbf{H}}}$, and the conditional distribution of GDM policy samples significantly resembles that of the expert policy. A peculiarity of optimal power control policies is that they tend to be probabilistic and involve multiple transmission modes. That is especially true for tx-rx pairs with less favorable channel conditions for which the policy randomization becomes more nuanced. In such cases, similar to time-sharing strategies, pairs that would otherwise generate considerable mutual interference adopt a policy-switching mechanism where they take turns to transmit at high power during periods of minimal interference (e.g., the top-left and bottom-right corners in the rightmost plot of Fig. 2). By alternating their transmissions, all tx-rx pairs satisfy their minimum ergodic rate requirements.
Strong test generalization evidenced in both figures notwithstanding, the GDM policy exhibits a small feasibility gap compared to the expert policy in Fig. 1. We attribute this gap primarily to the supervised training algorithm not accounting directly for the sensitivity of the constraints and the aforementioned policy-switching phenomenon. The feasibility gap and overall performance of the GDM policies can be further improved by incorporating the QoS requirements and additional variance constraints directly into the training loss and/or generative process.
@@ VI Conclusion
This work demonstrated that generative diffusion processes can imitate expert policies that sample from optimal solution distributions of stochastic network optimization problems. We employed a GNN to condition the generative process on a family of wireless network graphs. More broadly, we anticipate that our attempt at generative diffusion-based sampling of random graph signals will be of interest beyond resource optimization in wireless networks. We leave constraint-aware, unsupervised training of GDM policies across a wider range of network topologies as future research directions.